Through (-7,13), parallel to 4x+5y=12
Answer by Cromlix(4381)

4x+5y=12
Sort into y = mx + c
5y = -4x + 12
y = -4/5x + 12/5
Lines that are parallel have
the same gradient. = -4/5
Using y - b = m(x - a)
y - 13 = -4/5(x + 7)
y - 13 = -4/5x - 28/5
y = -4x/5 - 28/5 + 65/5
y = -4/5x + 37/5
OR
5y = -4x + 37
Hope this helps.
:-)